an object is displaced from a point A(0,0,0) to B(1m,1m,1m)under a force F=(yi+xj) Newton.The work done by this force in this process is

Grade:11

2 Answers

Arun
6 years ago
W = F. dr

W = (yi+xj) . (dxi+dyj)

dW = (ydx + xdy) 

dW = d(xy)

integrating both sides ,

W = xy from (0,0,0) to (1,1,1)

W = 1*1 - 0

W = 1 joule

Hope it helps

 

Regards

Arun(askIITians forum expert)

prasad
6 years ago

dW = (ydx + xdy)

dW = d(xy)

plz explain how do integrate ydx+xdy to get d(xy)